OpenStudy (anonymous):

please i need your help Application of counting Two cards are drawn at random from an ordinary deck of 52 cards (no jokers). Find the probability that the 2-card hand contains the given card. b.Nodeuces(2's) @tkhunny @satellite73

OpenStudy (anonymous):

\[\frac{48}{52}\times \frac{47}{51}\]